Sarah,Bob and Nellie are cousins. Sarah is half the age of Bob. In three years’ time Nellie will be twice as old as Sarah. Their ages at the moment add to 38 years. Determine the ages of Sarah, Bob and Nellie.
Answer by JulietG(1812)

B = 2S
N+3 = 2(S+3), which distributed becomes N+3=2S+6, or N=2S+3
S + B + N = 38
Take the known values of B and N from the top equations and put them into the bottom equation.
S + B + N = 38
S + (2S) + (2S+3) = 38
Add
5S + 3 = 38
Subtract 3 from each side
5S = 35
Divide each side by 5
S = 7
If Sarah is 7, then Bob is 14. Therefore, Nellie is 17 (38-21).
In three years, Sarah will be 10 and Nellie will be 20